OpenStudy (jazzyfa30):

Radioactive iodine is used to determine the health of the thyroid gland. It decays according to the equation where t is in days. Find the half-life of this substance. Round to the tenths place.For the half life: The half life is the solution (t) of the equation : a/2= ae ^ (-0.0856t)

OpenStudy (jazzyfa30):

@nincompoop

OpenStudy (anonymous):

typ there sorry

OpenStudy (anonymous):

solve \[\large e^{-.0865t}=.5\]

OpenStudy (jazzyfa30):

?????? idk what that is

OpenStudy (anonymous):

that's better takes only two steps \[-.0865t=\ln(.5)\] \[t=\frac{\ln(.5)}{-.0865}\]

OpenStudy (jazzyfa30):

ok can you walk me through the set up

OpenStudy (anonymous):

you have a typo in the problem i think Radioactive iodine is used to determine the health of the thyroid gland. It decays according to the equation MISSING EQUATION HERE where t is in days.

OpenStudy (anonymous):

i assumed the equation was \[\large A_0e^{-.0865t}\]

OpenStudy (anonymous):

is that correct?

OpenStudy (jazzyfa30):

i copied and pasted the entire Q

OpenStudy (anonymous):

yeah, but take a look i think when you copied and pasted it missed a piece

OpenStudy (jazzyfa30):

thats the way they had it on the Q see

OpenStudy (anonymous):

it doesn't matter, i can walk you through this, it is not too hard, and takes really only two steps to solve

OpenStudy (jazzyfa30):

ok

OpenStudy (anonymous):

at the very bottom you have this equation \[\large a/2=ae^{-.0865t}\] right?

OpenStudy (jazzyfa30):

yea

OpenStudy (anonymous):

the \(a\) is unnecessary, because the first step in solving is to divide both sides by \(a\) and get \[\large e^{-.0865t}=\frac{1}{2}=.5\]

OpenStudy (anonymous):

so far so good? we just cancelled the \(a\) from both sides of the equation

OpenStudy (jazzyfa30):

yea im following

OpenStudy (anonymous):

ok our goal to to get \(t\) by itself at the moment it is in the exponent the way to get it out of the exponent is to "take the natural log" of both sides, do you know what i am talking about? "no" is a fine answer, i am just wondering

OpenStudy (jazzyfa30):

im glad u stated that because no

OpenStudy (anonymous):

ok no problem

OpenStudy (anonymous):

the way to solve \(e^x=y\) for \(x\) is by writing in equivalent logarithmic form as \[x=\ln(y)\] you have two logs on your calculator, \(\log(x)\) which is base 10 and \(\ln(x)\) which is base \(e\) since \(e\) is our base, we clearly want to use \(\ln(x)\)

OpenStudy (anonymous):

in any case we will solve using wolfram, not a calculator so starting with \[\large e^{-.0865t}=.5\] we rewrite as \[-.0865t=\ln(.5)\]

OpenStudy (anonymous):

now the only think you need to do to get \(t\) by itself is to divide both sides by \(-.0865\) final answer is \[\large t=-\frac{\ln(.5)}{.0865}\]

OpenStudy (anonymous):

looks like about 8 http://www.wolframalpha.com/input/?i=-ln%28.5%29%2F.0865

OpenStudy (anonymous):

not sure how clear that was, but i hope you see it was really only two steps

OpenStudy (jazzyfa30):

no thats incorrect

OpenStudy (anonymous):

it does say "round to the tenths place", right?

OpenStudy (anonymous):

ah damn, i put it \(.0865\) instead of \(.0856\)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-ln%28.5%29%2F.0856 try \(8.1\)

OpenStudy (jazzyfa30):

nope